### v1

@@ IV-C LSC’24: Performance Analysis
In this section, we take a closer look at the performance of systems in the LSC’24. Novice tasks were excluded from this analysis due to wide variability in system performance. In each figure, the teams are ordered by their overall score in the respective task type (KIS, QA, Ad-hoc), which can be seen in the leaderboard in Figure IV-C. We follow the same structure as in Section IV, focusing on the correctness of submissions, time to the first correct submission, and overall scores.
@@ IV-C1 Correctness of Submissions
We observed a direct correlation between the number of correct submissions and overall scores in KIS and QA tasks, as shown in Figure IV-C1. Teams that submitted more correct answers tended to achieve higher scores. Lower-scoring teams often made multiple incorrect attempts, which reduced their overall performance. In ad-hoc tasks, since the number of groundtruth images varied widely across queries, the number of correct submissions was less indicative of overall performance as the scores were normalized per task. Therefore, we focused on precision and recall per task to evaluate system performance in ad-hoc tasks, as shown in Figure IV-C1. In this graph, we observed a stronger correlation between recall and overall scores, compared to precision. Even though penalties were applied for incorrect submissions, recall played a more significant role in determining the final score as described in the evaluation methodology. Nevertheless, the top-performing teams still maintained high precision, indicating a balanced approach.
@@ IV-C3 Score Distribution
In order to understand more about the performance of teams in LSC’24, we analysed the distribution of scores across each individual tasks in Figure IV-C3. The heatmap reveals clear distinctions in performance, highlighting top teams, task-specific challenges, and consistency across different retrieval methods.
@@
Upon closer inspection, KIS04 presents an anomaly, as all teams scored zero, which is highly unusual for KIS tasks. The query hints included: ‘It was as if the meeting took place on the Starship Enterprise. I remember the wall had yellow lighted shapes and designs, with a yellow lighting on the ceiling. There were 3 or 4 people in the meeting room with me. It was a formal boardroom. Afterwards, I got a taxi back to my hotel in Bangkok.’ This query was conceptually vague, relying on subjective interpretation rather than concrete object-based details. The metaphorical reference to the Starship Enterprise may have confused both users and retrieval models, making it difficult to identify relevant images. By the time specific visual details were provided (e.g., yellow lighting, boardroom, people in the room), users may have already formed a completely different mental image, leading to misinterpretations. This anomaly underscores the importance of clear, structured query design to ensure that retrieval tasks remain solvable.
@@
Ad-hoc tasks posed the greatest challenge, with notable lower scores across teams. The average scores for these tasks were under 23, while median scores remained below 14. While LIFEXPLORE1, VISIONE1, and SnapSeek1 maintained competitive results, many teams struggled to achieve meaningful retrieval results, often scoring near zero in multiple tasks. One notable case was LSC24-AD05, where PraK2 achieved 93, while PraK1 scored 63, creating a large disparity that made other teams appear weaker due to normalization effects in the pooled evaluation metric. This suggests that some ad-hoc queries may be highly system-dependent, favoring specific retrieval models over others.
Lower-performing teams should refine their strategies for open-ended ad-hoc tasks to improve future performance, as these exhibited the greatest variability and lowest overall scores. Even the top teams could benefit from improving consistency across different retrieval types, ensuring robust performance not only in structured KIS queries but also in more semantically complex ad-hoc and QA tasks. Addressing these challenges will be crucial for advancing lifelog retrieval systems and improving their adaptability to diverse search scenarios.
@@ IV-D Insights on Techniques and System Performance
This section summarises key insights regarding how the retrieval techniques employed by different teams influenced their overall system performance across the LSC tasks. Even though the performance of systems varied widely, the following trends can be observed:
@@
Ultimately, the most significant insight from this analysis is that no single retrieval technique or design choice conclusively predicted system performance. Instead, successful lifelog retrieval systems seem to depend upon careful integration, tuning, and balanced design choices rather than any isolated retrieval method or technology. Additionally, deploying multiple instances of the same system in expert-track evaluations demands careful management of user familiarity. Such multi-instance designs might better suit the novice track in future LSC editions, where variability due to user experience can be more meaningfully assessed.
@@ V-B System Performance and Key Observations
Scores in KIS and QA tasks improved from 2022 to 2023 but became more variable in 2024 due to the introduction of multiple system instances with varied configurations. Correct submissions were closely correlated with the number of available groundtruth images, with 2023’s KIS tasks benefitting from a higher count, while 2024’s reduced number of groundtruth images increased difficulty.
Performance varied notably across task types. Ad-hoc tasks remained the most challenging, displaying wide variance in team performance on precision and recall. Systems such as LIFEXPLORE, SnapSeek, and VISIONE demonstrated consistently high scores, while others exhibited fluctuations due to differences in system configurations. Additionally, LLM-integrated systems like MemoriEase, Memento 4.0, and Voxento-Pro excelled in QA tasks, showcasing the effectiveness of conversational search and RAG techniques in refining retrieval precision. The failure of KIS04, however, underscored how abstract or ambiguous prompts, while arguably closer to natural memory recall, can be particularly difficult for current retrieval systems. This highlights the ongoing challenge in KIS task design: while realistic, memory-like queries are desirable, they must still be feasible for systems operating without prior user context. This case demonstrates the importance of calibrating ambiguity and providing enough grounding to enable meaningful evaluation. However, as the LSC enters its eighth year, it is appropriate for tasks to become more challenging to test the evolving capabilities of state-of-the-art systems. Realistic, memory-like queries—often imprecise or fuzzy—better reflect real-world information needs in lifelog retrieval.
@@ V-C Limitations
This paper provides an overview of system performance in LSC 2024, but several limitations must be acknowledged. First, the analysis is based on a dataset collected from a single individual, which limits its demographic, cultural, and behavioural representativeness. While it enables longitudinal analysis and consistent annotation, the findings may not generalise to more diverse lifelogging contexts. However, we would like to note that the lifelogger traveled to many countries across multiple continents, thereby capturing a wide variety of cultural contexts. The unique strength of the dataset lies in the continuous, long-term data collection, offering rare insights into behavioral patterns over time.
Second, we excluded novice systems from quantitative analysis due to their inconsistency and incompleteness, which may limit insights into first-time usability. Performance variations among systems with shared cores (e.g., SnapSeek variants) further reflect the influence of user factors outside the scope of this analysis.
Third, while all systems within a given LSC edition are evaluated on the same set of tasks under identical conditions, ensuring fairness within each year, comparisons across years are more nuanced. Variability in task difficulty, especially in subjective or open-ended tasks, may influence performance trends. Although we mitigate this by averaging scores across multiple tasks and reporting per-task-type performance, a fully consistent difficulty calibration across years remains an open challenge. Achieving this would likely require re-running historical systems on current-year tasks, ideally with the same users and interfaces-something not currently feasible in the live competitive format.
Additionally, Gini was used to reflect temporal imbalance due to its interpretability and simplicity. We acknowledge it is a simplification, and alternatives such as entropy or burstiness could offer richer descriptions of temporal structure. However, given space constraints and our focus on system behavior rather than metric development, we leave deeper exploration to future work.
@@ V-D Future Directions
The findings from LSC’24 suggest several areas for future development in lifelog retrieval systems:
  • Refinement of Conversational Search: While LLMs have improved interactive retrieval, there remains room for enhancing context-awareness and response accuracy. Conversational search may also help users refine ambiguous queries iteratively, which complements foundational retrieval efforts. Future iterations should explore fine-tuning models on lifelog-specific datasets.
  • Optimized Ad-hoc Query Handling: Addressing performance disparities in Ad-hoc retrieval requires improved search strategies, including better query expansion, relevance feedback mechanisms, and dynamic result ranking.
  • Enhanced Personalization and Adaptability: Systems could integrate user-specific preferences and search behaviors to provide more tailored retrieval experiences. This could involve hybrid approaches that combine embeddings, knowledge graphs, and user feedback loops.
  • Improved Evaluation Metrics: Given the increasing complexity of lifelog retrieval tasks, more nuanced evaluation metrics that account for query complexity, retrieval diversity, and semantic similarity could provide better performance insights.
  • Multimodal Fusion and Advanced Interfaces: Integrating text, voice, and image-based inputs can enable users to express their queries more intuitively. Virtual and augmented reality interfaces may further enhance user interaction by providing immersive environments where users can “relive” events through reconstructed lifelog data. However, practical challenges remain in adopting these interfaces widely.
  • Event Summarization and Automated Narratives: As lifelogs capture increasingly granular details, summarization tools that generate coherent narratives from raw data will become essential. These tools could help users quickly review their daily activities, significant milestones, or recurring patterns, improving accessibility and usability.
  • Collaborative Search Environments: Collaboration features, such as shared lifelog workspaces and annotation tools, can support joint exploration of lifelog data for purposes such as health monitoring, caregiving, or memory sharing. Developing secure and user-friendly collaborative search environments will enable users to co-navigate their lifelog collections with trusted individuals while addressing privacy and security concerns.
Especially, given our query difficulty analysis in Section IV-B, future work could explore providing optional guidance to lifeloggers during query formulation. Some initial pointers could include:
  • Queries should aim to reflect realistic memory recall but avoid phrasing that is too vague or abstract, especially if no strong visual cues are present. Instead, prompts should strike a balance between authenticity and retrievability
@@
  • Also for Ad-hoc tasks, one promising direction is to move beyond scoring individual images and instead evaluate system outputs at the level of temporal segments. Many Ad-hoc queries implicitly refer to extended activities (e.g., ‘using chopsticks at a meal’, ‘eating fast food’), and this approach would allow evaluation metrics to reflect how well systems retrieve relevant episodes, not just isolated images. The current image-level scoring allows for a potential shortcut: once a system retrieves a single image from a long groundtruth event, it can submit the surrounding frames to accumulate high scores, even if their overall retrieval strategy lacks precision or coverage. Segment-based scoring would reduce the impact of such strategies by emphasising temporal coherence and meaningful coverage. Although this is not currently feasible due to the lack of a shared segmentation standard, we see clear motivation for shifting from image-level to segment-level metrics. We present this not as an immediate solution, but as a promising direction for future editions of the challenge. Moreover, this shift would also necessitate the development of new evaluation protocols, including metrics for temporal overlap and segment-level recall and precision.
@@
Another important direction is addressing dataset scarcity without compromising participant privacy. Synthetic data generation (e.g., [74]) and simulation-based lifelogging offer privacy-preserving alternatives. In parallel, repurposing existing egocentric video datasets, such as EPIC-Kitchens [14], Ego4D [21], or the recently released EgoLife [93] and CASTLE [62], provides a promising route for evaluating lifelog-style retrieval systems in naturalistic settings. While these datasets are not full lifelogs, task-specific annotations can adapt them for lifelogging benchmarks. We encourage future work to explore shared annotation efforts on such datasets and to develop protocols for synthetic, multi-perspective lifelog data collection.
@@ V-E Final Thoughts
Although the LSC takes place in a constrained, competitive environment, many of its task formulations closely mirror real-world lifelogging scenarios—such as recalling past events, locating personal media, or answering context-rich questions. This alignment suggests that the findings from LSC submissions are not only useful for benchmarking system performance, but also transferable to the development of practical, user-facing lifelog retrieval tools.
@@
Allie Tran: Conceptualization, Formal analysis, Investigation, Writing – Original Draft, Visualization, Writing – Review & Editing

### v2

@@ IV-C LSC’24: Performance Analysis
In this section, we take a closer look at the performance of systems in the LSC’24. Novice tasks were excluded from this analysis due to wide variability in system performance. In each figure, the teams are ordered by their overall score in the respective task type (KIS, QA, Ad-hoc), which can be seen in the leaderboard in Figure IV-C. We follow the same structure as in Section IV, focusing on the correctness of submissions, time to the first correct submission, and overall scores. This follows the same sequence as before, allowing direct comparison between cross-year trends and 2024-specific results.
@@ IV-C1 Correctness of Submissions
We observed a direct correlation between the number of correct submissions and overall scores in KIS and QA tasks, as shown in Figure IV-C1. Teams that submitted more correct answers tended to achieve higher scores. Lower-scoring teams often made multiple incorrect attempts, which reduced their overall performance. In ad-hoc tasks, since the number of groundtruth images varied widely across queries, the number of correct submissions was less indicative of overall performance as the scores were normalised per task. Therefore, we focused on precision and recall per task to evaluate system performance in ad-hoc tasks, as shown in Figure IV-C1. In this graph, we observed a stronger correlation between recall and overall scores, compared to precision. Even though penalties were applied for incorrect submissions, recall played a more significant role in determining the final score as described in the evaluation methodology. Nevertheless, the top-performing teams still maintained high precision, indicating a balanced approach. Next, we examine whether timing can help explain the variation observed in correctness and precision.
@@ IV-C3 Score Distribution
In order to understand more about the performance of teams in LSC’24, we analysed the distribution of scores across each individual task in Figure IV-C3. The heatmap reveals clear distinctions in performance, highlighting top teams, task-specific challenges, and consistency across different retrieval methods.
@@
Upon closer inspection, KIS04 presents an anomaly, as all teams scored zero, which is highly unusual for KIS tasks. The query hints included: ‘It was as if the meeting took place on the Starship Enterprise. I remember the wall had yellow lighted shapes and designs, with a yellow lighting on the ceiling. There were 3 or 4 people in the meeting room with me. It was a formal boardroom. Afterwards, I got a taxi back to my hotel in Bangkok.’ This query was conceptually vague, relying on subjective interpretation rather than concrete object-based details. The metaphorical reference to the Starship Enterprise may have confused both users and retrieval models, making it difficult to identify relevant images. By the time specific visual details were provided (e.g., yellow lighting, boardroom, people in the room), users may have already formed a completely different mental image, leading to misinterpretations. This anomaly underscores the importance of clear, structured query design to ensure that retrieval tasks remain solvable. This also illustrates a broader limitation of current multimodal retrieval: while embedding models align literal or concrete descriptions well, they still struggle with figurative, cultural, or metaphorical language. Bridging this ‘semantic gap’ between abstract phrasing and visual grounding remains an open research challenge, demanding hybrid models that integrate world knowledge and commonsense reasoning with visual grounding.
@@
Ad-hoc tasks posed the greatest challenge, with notably lower scores across teams. The average scores for these tasks were under 23, while the median scores remained below 14. While LIFEXPLORE1, VISIONE1, and SnapSeek1 maintained competitive results, many teams struggled to achieve meaningful retrieval results, often scoring near zero in multiple tasks. One notable case was LSC24-AD05, where PraK2 achieved 93, while PraK1 scored 63, creating a large disparity that made other teams appear weaker due to normalisation effects in the pooled evaluation metric. This suggests that some ad-hoc queries may be highly system-dependent, favouring specific retrieval models over others.
Lower-performing teams should refine their strategies for open-ended ad-hoc tasks to improve future performance, as these exhibited the greatest variability and lowest overall scores. Even the top teams could benefit from improving consistency across different retrieval types, ensuring robust performance not only in structured KIS queries but also in more semantically complex ad-hoc and QA tasks. Addressing these challenges will be crucial for advancing lifelog retrieval systems and improving their adaptability to diverse search scenarios. In sum, LSC’24 mirrors the cross-year story: high KIS solvability, volatile QA difficulty, and ad-hoc sensitivity to instance/user effects and query design.
@@ IV-D Insights on Techniques and System Performance
Finally, we synthesise the above results into technique-level insights, tying method choices back to the correctness, timing, and score patterns. This section summarises key insights regarding how the retrieval techniques employed by different teams influenced their overall system performance across the LSC tasks. Even though the performance of systems varied widely, the following trends can be observed:
@@
Ultimately, the most significant insight from this analysis is that no single retrieval technique or design choice conclusively predicted system performance. Instead, successful lifelog retrieval systems seem to depend upon careful integration, tuning, and balanced design choices rather than any isolated retrieval method or technology. Additionally, deploying multiple instances of the same system in expert-track evaluations demands careful management of user familiarity. Such multi-instance designs might better suit the novice track in future LSC editions, where variability due to user experience can be more meaningfully assessed. Taken together, these points unify the metric trends and query-difficulty results, concluding the section’s analysis.
@@ V-B System Performance and Key Observations
Scores in KIS and QA tasks improved from 2022 to 2023 but became more variable in 2024 due to the introduction of multiple system instances with varied users or configurations. Correct submissions were closely correlated with the number of available groundtruth images, with 2023’s KIS tasks benefiting from a higher count, while 2024’s reduced number of groundtruth images increased difficulty.
Performance varied notably across task types. Ad-hoc tasks remained the most challenging, displaying wide variance in team performance on precision and recall. Systems such as LIFEXPLORE, SnapSeek, and VISIONE demonstrated consistently high scores, while others exhibited fluctuations due to differences in instances. Additionally, LLM-integrated systems like MemoriEase, Memento 4.0, and Voxento-Pro excelled in QA tasks, showcasing the effectiveness of conversational search and RAG techniques in refining retrieval precision. The failure of KIS04, however, underscored how abstract or ambiguous prompts, while arguably closer to natural memory recall, can be particularly difficult for current retrieval systems. This highlights the ongoing challenge in KIS task design: while realistic, memory-like queries are desirable, they must still be feasible for systems operating without prior user context. This case demonstrates the importance of calibrating ambiguity and providing enough grounding to enable meaningful evaluation. However, as the LSC enters its eighth year, it is appropriate for tasks to become more challenging to test the evolving capabilities of state-of-the-art systems. Realistic, memory-like queries—often imprecise or fuzzy—better reflect real-world information needs in lifelog retrieval.
@@ V-C Limitations
This paper provides an overview of system performance in LSC 2024, but several limitations must be acknowledged. First, while privacy and ethics are central to lifelogging in general, this work focuses on retrieval performance under an already ethically approved dataset.
Second, the analysis is based on a dataset collected from a single individual, which limits its demographic, cultural, and behavioural representativeness. While it enables longitudinal analysis and consistent annotation, the findings may not generalise to more diverse lifelogging contexts. However, we would like to note that the lifelogger travelled to many countries across multiple continents, thereby capturing a wide variety of cultural contexts. The unique strength of the dataset lies in the continuous, long-term data collection, offering rare insights into behavioural patterns over time.
Third, we excluded novice systems from quantitative analysis due to their inconsistency and incompleteness, which may limit insights into first-time usability. Performance variations among systems with shared cores (e.g., SnapSeek variants) further reflect the influence of user factors outside the scope of this analysis.
Moreover, while all systems within a given LSC edition are evaluated on the same set of tasks under identical conditions, ensuring fairness within each year, comparisons across years are more nuanced. Variability in task difficulty, especially in subjective or open-ended tasks, may influence performance trends. Although we mitigate this by averaging scores across multiple tasks and reporting per-task-type performance, a fully consistent difficulty calibration across years remains an open challenge. Achieving this would likely require re-running historical systems on current-year tasks, ideally with the same users and interfaces-something not currently feasible in the live competitive format.
Additionally, Gini was used to reflect temporal imbalance due to its interpretability and simplicity. We acknowledge it is a simplification, and alternatives such as entropy or burstiness could offer richer descriptions of temporal structure. However, given space constraints and our focus on system behaviour rather than metric development, we leave deeper exploration to future work.
@@ V-D Future Directions
Building on our longitudinal analysis and query difficulty findings, we outline future research directions that address observed system trends, evaluation limitations, and emerging interaction challenges in lifelog retrieval.
  • Refinement of Conversational Search: While LLMs have improved interactive retrieval, there remains room for enhancing temporal and situational context awareness and improving response accuracy. Conversational search may also support iterative refinement of ambiguous queries through follow-up prompts or clarifications, complementing foundational retrieval efforts. Future iterations require investigation of fine-tuning models on lifelog-specific datasets to better capture personal, temporal, and activity-related semantics.
  • Optimised Ad-hoc Query Handling: Addressing performance disparities in Ad-hoc retrieval requires improved search strategies, specifically through automated query expansion (using generative AI, etc.) and interactive negative relevance feedback that allows users to quickly filter out visually similar but contextually irrelevant false positives.
  • Improved Evaluation Metrics: Given the increasing complexity of lifelog retrieval tasks, more nuanced evaluation metrics that account for query complexity, retrieval diversity, and the cognitive or interaction cost could provide better performance insights. Specifically, we suggest adopting graded relevance and Interaction-Cost-to-Success to measure search efficiency alongside existing metrics.
  • Controlled Studies with Novice Users While expert performance provides a controlled view of system capabilities, future work should systematically examine user variability under realistic conditions. Controlled studies could recruit both novices and semi-experienced users to assess how familiarity, interface complexity, and retrieval transparency affect task efficiency and satisfaction. Mixed-method evaluations—combining task analytics with think-aloud or post-task reflection—would enable quantifying learning effects and uncovering usability barriers specific to first-time lifelog users.
Especially given our query difficulty analysis in Section IV-B, future work could explore optional guidance for lifeloggers during query formulation, with the aim of balancing authenticity of memory recall and retrievability. Some initial pointers could include:
  • Queries should aim to reflect realistic memory recall but avoid phrasing that is too vague or abstract, especially if no strong visual cues are present. Instead, prompts should strike a balance between authenticity and retrievability.
@@
  • Also for Ad-hoc tasks, one promising direction is to move beyond scoring individual images and instead evaluate system outputs at the level of temporal segments. Many Ad-hoc queries implicitly refer to extended activities (e.g., ‘using chopsticks at a meal’, ‘eating fast food’), and this approach would allow evaluation metrics to reflect how well systems retrieve relevant episodes, not just isolated images. The current image-level scoring allows for a potential shortcut: once a system retrieves a single image from a long groundtruth event, it can submit the surrounding frames to accumulate high scores, even if their overall retrieval strategy lacks precision or coverage. Segment-based scoring would reduce the impact of such strategies by emphasising temporal coherence and meaningful coverage. Moreover, adopting a segment-based approach better reflects realistic retrieval behaviour, as users tend to search for moments or episodes rather than individual frames. Segment-level metrics further assess a system’s ability to maintain temporal continuity across related scenes. Although this is not currently feasible due to the lack of a shared segmentation standard, we see clear motivation for shifting from image-level to segment-level metrics. We view the transition from image-level to segment-level evaluation not as an immediate solution, but as a longer-term research agenda requiring shared segmentation standards and new temporal evaluation metrics.
@@
Finally, addressing dataset scarcity while preserving participant privacy remains a foundational challenge for lifelog research. Synthetic data generation (e.g., [75]) and simulation-based lifelogging offer privacy-preserving alternatives. In parallel, repurposing existing egocentric video datasets, such as EPIC-Kitchens [14], Ego4D [21], or the recently released EgoLife [94] and CASTLE [63], provides a promising route for evaluating lifelog-style retrieval systems in naturalistic settings. While these datasets are not full lifelogs, task-specific annotations can adapt them for lifelogging benchmarks. We encourage future work to explore shared annotation efforts on such datasets and to develop protocols for synthetic, multi-perspective lifelog data collection.
@@ V-E Final Thoughts
Although the LSC takes place in a constrained, competitive environment, many of its task formulations closely mirror real-world lifelogging scenarios—such as recalling past events, locating personal media, or answering context-rich questions. This alignment suggests that the findings from LSC submissions are not only useful for benchmarking system performance but also transferable to the development of practical, user-facing lifelog retrieval tools.
@@
Allie Tran: Conceptualisation, Formal analysis, Investigation, Writing – Original Draft, Visualisation, Writing – Review & Editing